# New York, US

### 05 Jan 2024 v1

@@ I Introduction
Bloom Filters [3] are a time-tested, space-efficient data structure for identifying duplicate items within an input sequence, and have been applied in an extremely broad set of computing applications, ranging from packet processing and forwarding/routing in P2P networks [25], to cache summarization and cache filtering in CDNs [16], network monitoring [12], web caching [1], network security enhancement [8], database search [29], navigation tasks in mobile devices [11], data synchronization [15], and even Biometric authentication [18]. The Bloom Filter (BF) is attractive because it errs only in the False Positive direction: an arriving input, which we call a message, can be incorrectly identified as a repeat of a previous message and it never errs in the false negative direction, i.e., never misclassifying a repeat message as new.
There are several variants of “back-of-the-envelope” analyses [^1] Later in the paper, we look at different variants. that give the False Positive rate $f$ of an $n+1$st message after $n$ messages have been inserted, e.g., one common variant yields $f=(1-(1-1/M)^{kn})^{k}$, where $M$ is the memory (number of bits in the BF) and $k$ is the number of hash functions, each mapping the message to one of the $M$ bits. This analysis, while useful to demonstrate efficacy of the BF, does not measure False Positives in a way that best serves practical network applications for two reasons:
Observability: First, consider the addition of an $n$th element. If the application maintains a count, $b$, of the number of bits set, the next new arrival being classified as a repeat (a False Positive) can be directly computed as $(b/M)^{k}$: the earlier analysis disregards the ability to inspect current BF state.
@@
A class of applications that would benefit from a measure of False Positive rate that takes into account both observability and average rate are those that employ what we call a recycling Bloom Filter (RBF) [16, 26, 19]. These applications utilize BF’s across sufficiently long timescales, where new messages are continuously arriving, but are interspersed with repeats of previous messages. Since newly arriving messages almost always set bits in the BF, over time the number of bits will become excessive, driving up the False Positive rate, such that some bits must be cleared to keep False Positive rates at a reasonable level. While there are BF variants that do support deletion, their implementation introduces substantial overhead or requires knowledge of the specific elements destined for deletion [22, 13, 9, 2]. Such overhead and complexity is not suited for network applications where message arrivals have a “fading popularity” character, with repeat probability decreasing over time. For applications with this type of message arrival process, one can utilize the standard BF that maintains some measure of BF fullness, such as the number of messages inserted, or the number of bits set in the BF. When this measure exceeds a threshold, the BF recycles: all bits are cleared, and the filling process begins anew (we say a new BF cycle is initiated). This process can continue indefinitely, providing a simple and effective means for maintaining a low maximal (and average) False Positive rate of arrivals.
In this paper, we develop recursive, quickly computable (on today’s conventional hardware) analytical models that accurately capture the average False Positive rates of recycling BFs (RBFs). We show that using such an active measure and gauging the recycle to maintain an average (as opposed to maximum) False Positive rate can offer a $30$% or more increase in amount of messages stored.
@@
  • We verify our models using discrete-event-driven simulation in §V and using our models, contrast performance of the RBF variants as functions of memory size and desired False Positive rates.
@@
Our findings show that the variant that recycles based on the number of bits set, while most complex to model to find the right parameters, is the best performing variant in that it can maintain a given False Positive rate while packing in the largest expected number of messages prior to recycling. In contrast, the variants that recycle based on message counts to achieve an average False Positive rate are easier to model and achieve only a slightly smaller expected number of messages prior to recycling. These average-case variants store around 30% more messages in expectation than the commonly used worst case variant. We also find that implementing a two-phase variant will closely match the one-phase as the False Positive rate to be met decreases. These results are important in that they point to users of BFs (and specifically RBFs) likely being overly conservative for their needs when using previous measures of False Positive.
@@ II Background
$f_{w}$ False Positive rate using worst-case threshold
$f_{o}$ False Positive rate using fixed $N$ oracle lower bound
$f_{a}$ f.p. rate using fixed $N$ bit-setting msgs lower bound
one-phase RBF
$f_{\sigma}^{2}$ False Positive rate using $\sigma$ bound, two-phase RBF
@@ II-A Bloom Filters
When a message arrives, the bits to which it hashes are checked before being set. If they are all already set, the message is assumed to be a repeat of a prior message, since a repeat message’s bits will always already be set by the prior arrival of the message to the BF. In the rare instance that a new message’s $k$ hashes point to bits already set, it is incorrectly classified as a repeat, i.e., a False Positive.
@@
Conventional implementations of hash functions permits several of the $k$ hash functions mapping to collide and hash to the same bit (i.e., a pseudo-random sampling with replacement). A slightly more sophisticated hashing scheme can ensure the $k$ hashes for a message are distinct from one another (i.e., a pseudo-random sampling without replacement). We respectively refer to these two Bloom Filter variants as colliding and non-colliding, and we analyze both types. While we show how to incorporate these variants into our models, for values of $M$ and $k$ used in practice (i.e., $M>1,000$), the variant of hash function (colliding vs. non-colliding) has an unnoticeable impact on False Positive rate (collisions are sufficiently rare in the colliding variant). Note that for either variant of hash functions, the independence of bits assigned to differing messages still holds. Hence, our results are generally presented using the colliding variant.
@@ II-B Traditional False Positive Analysis
where, for a given $M$ and $n$, $k\approx\frac{M}{n}\ln(2)$ will minimize the corresponding False Positive rate. Wikipedia provides a succinct derivation of this simple formula [27], as well as a formula for non-colliding hash functions (in terms of Stirling numbers). From (1), the False Positive rate approaches $1$ as the number of messages $n$ increases: an undesirable property for applications with no limit on the number of potential message insertions. We call using such a bound in practice to decide when to recycle the BF a worst case bound (hence $f_{w}$) because the recycling occurs as soon as a particular message’s False Positive rate exceeds $f_{w}$ (while the average over messages inserted will be lower).
In practice, these applications allow the BF to fill with messages until a predetermined maximum False Positive rate is reached, after which time it is “recycled”, where all bits are reset to 00 and subsequent message arrivals again fill the BF. We refer to the lifetime of a BF between recycling events as a cycle of the BF: the cycle ends when the BF recycles.
@@ II-C Two-Phased Bloom Filter
A recycled BF (RBF), upon recycling, clears all memory with respect to previous arrivals. To mitigate the effects of this memory loss, a BF variant called a Two-Phased Bloom Filter has been used in various network applications, e.g., [16, 26].
The two-phase Bloom Filter can be thought of as two arrays with $M$ bits each, indexed $10,...,M-$ 1, along with $k$ independent hash functions that are shared across both arrays. Typically, $M=\frac{M^{\prime}}{2}$, where $M^{\prime}$ is what the application’s memory constraint would be for the standard (one phase) Bloom Filter.
The two-phase variant operates much like the standard recycling variant, with the addition that at any point in time, one Filter array is designated active and the other frozen. Incoming messages are hashed and added to the active filter. When the active filter is considered to be “full enough” (e.g., the number of bits set exceeds or the number of messages stored exceeds some threshold), instead of clearing the bits of the active filter, we clear the bits of the frozen filter. The roles of the active and frozen filter are then switched; incoming messages are added to the previously frozen (now active) filter, and the status of the previously active (now frozen) filter remains unaffected. This process repeats indefinitely.
In the two-phase variant, an arriving message is hashed and its bits are compared separately within both the active and frozen filters. A match with either filter is interpreted as the message having been received previously (and having set the corresponding bits in the respective filter).
@@ II-D BF Thresholding Approaches
There are two basic approaches to deciding when to recycle a BF. One can count the number of bits $b$ that are set in the BF and recycle when $b$ exceeds a threshold $\sigma$. We call this approach a $\sigma$-bounded approach. Alternatively, one can bound the number of bit-setting messages, $N$, that can be admitted into the BF before recycling. [^2] Messages that don’t set bits are generally not counted because the user must assume they are repeat messages, and since repeat messages never set bits, the False Positive rate of the BF does not increase due to their repeated arrival. We call this approach an $N$-bounded approach. We note that one complication with an $N$-bounded approach is that a user cannot distinguish when a message that arrives without setting any bits is a repeat message, or is a False Positive, i.e., a new message that hashed to bits that were already set in the BF. Note that formulas for an $N$-bounded approach, such as (1) count new messages regardless of whether or not new bits are set.
@@ II-E Averaging Variants
A BF is only of use when the message arrival process consists both of new and repeat messages. We wish to clarify a subtle point about what we are averaging over. When a message first arrives, if it sets additional bits, it is identified as new, and if not, it is classified as a repeat. Consider the $i$th new message that arrives, and let $\eta(i)$ represent the number of times the message arrives within a cycle, and let $F_{i}$ be an indicator that equals 1 if message $i$ upon first arrival is a False Positive. We can count these $\eta(i)$ arrivals in three different ways.
@@
Which variant to use depends on the needs of the underlying application. For instance, a False Positive in a caching application that uses the BF to determine cache hits for requests would make the wrong assessment for an initial arrival of a request, but the mistake would be detected upon attempting to fetch the item from the cache, so the item would be cached and subsequent requests would not be false positives. Hence, count-first is the appropriate measure. Alternatively, a mechanism that wishes to count unique messages would have a single miscount for each message initially classified incorrectly as having been previously received, hence count-instance is the appropriate measure. Finally, in a setting where the BF is used to drop repeat requests, and where a penalty is paid per dropping of an unsatisfied request, count-instance would be the appropriate measure.
We can prove that count-instance is the most stringent (largest) of False Positive rates of the three versions333count-first is straightforwardly smaller count-each.. The proof showing that count-instance results in the largest False Positive rate is omitted here but is included in an extended version [OurExtendedVersion] of this publication. This observation yields two benefits. First, if we compute the false positive rate for count-instance, we have upper bounds on the rates of the count-first and count-each variants. [^4] In fact, our analysis of count-instance is itself an upper-bound in an RBF for a subtle reason that we address in the extended version. Second, count-instance is not affected by underlying distribution of arriving messages: how often a particular message repeats is of no import, since we only decide its contribution to False Positive rate based on its first arrival, and any subsequent arrivals do not alter the bits in the BF, making its analysis identical regardless of this underlying distribution, whereas the other variants do depend on the underlying distribution.
@@ III $N$-bounded average False Positive rate
First, let us consider an oracle user, who inserts new messages into the BF and “knows” that even when a message sets no additional bits in the BF that it is new. [^5] This is a hypothetical user because the user who knows this has no need for a BF. This oracle would insert exactly $N$ new messages, and the resulting average False Positive rate can be computed as a simple renewal process, yielding an oracle $N$-bounded average False Positive rate of:
@@
In contrast, a “real” user would assume that a new message that sets no bits is instead a repeat message, and would not include such a message in their count of $N$ messages. For this real user, we let $i$ iterate over the number of new messages that set bits, and define r.v. $R_{i}$ to equal the number of new messages that do not set bits and arrive between the $i-1$st and $i$th new messages that do set bits. Then for a single cycle of the RBF, the False Positive rate is $\sum_{i=1}^{N}R_{i}/\sum_{i=1}^{N}(R_{i}+1)$.
Note that for the messages that comprise a given $R_{i}$, the number of bits set in the BF remains constant (none of the $R_{i}$ messages are changing the number of bits set), so each such message has an equal false probability likelihood, which we call $p_{i}$, of not setting additional bits. Note it follows that $P(R_{i}>j)={p_{i}}^{j+1}$, such that $E[R_{i}]=p_{i}(1-p_{i})^{-1}$. By renewal theory, we have the False Positive rate is thus equal to
@@
Finally, we claim that $p_{i}$ is lower-bounded by $P(F_{i}=1)$. This follows from the fact that the $i$th new message that set bits is actually the $j$th arriving message with $j\geq i$, so $p_{i}=P(F_{j}=1)$ for some $j\geq i$. The lower bound then follows from the fact that $P(F_{i}=1)$ is increasing in $i$ (False Positive rates increase as more messages arrive), and that the worst-case analyses’ value for $N$ is over all messages, including those that set no bits. Hence, $p_{i}=P(F_{j}=1)\geq P(F_{i}=1)$ for $j\geq i$. Since we are underestimating each $p_{i}$, we are underestimating the likelihood of new arrivals not setting bits, so we are underestimating False Positive rates. Note that for low False Positive rates that are used in practice, $E[R_{i}]$ will tend to be very small, such that the bound is expected to be tight.
It will be useful for our purposes to define, for $f_{i}=P(F_{j}=1)$ directly applying (1), we have an actual $N$-bounded False Positive rate of
@@ IV $\sigma$-bounded average False Positive rate
Figure 1: Partial Markov Chain diagram for Recycling Bloom Filter, showing the “forwards” transition behavior, for $k=2$. States represent the number of RBF bits set to $1$ and $\tau_{k}{i,j}$ is the probability of transitioning from state $i$ to state $j$.
As a starting point towards developing a recursive analytical model that accurately captures the average False Positive rates for the $\sigma$-bounded model, we consider a process where messages sequentially arrive at the RBF. Once a message arrives, the $k$ hash functions are applied to it and the corresponding bits are set to $1$ (or remain $1$ if already set). In a departure from the traditional method of Bloom Filter recycling, we do not reset the RBF after $n$ unique message insertions. Instead, we have a predetermined bit capacity parameter $\sigma\leq M$. After the insertion of a message, let $b$ be the number of bits set to 1 in the RBF. We reset the RBF when $b>\sigma$. Our recursive Markov Model will allow us to determine the long-term average False Positive rate across all new arrivals for a given value of $\sigma$ (another departure from the traditional False Positive analysis, which only considers the “worst case” False Positive rate for the last arrival of a RBF reset cycle).

### 03 Feb 2024 v2

@@ I Introduction
Bloom Filters [3] are a time-tested, space-efficient data structure for identifying duplicate items within an input sequence, and have been applied in an extremely broad set of computing applications, including packet processing and forwarding/routing in P2P networks [19], cache summarization and cache filtering in CDNs [13], network monitoring [9], data synchronization [12], and even Biometric authentication [15]. The Bloom Filter (BF) is attractive because it errs only in the False Positive direction (an arriving input, which we call a message, can be incorrectly identified as a repeat of a previous message) and never the False Negative direction (a repeat message is never incorrectly classified as new).
There are several variants of “back-of-the-envelope” analyses [^1] Later in the paper, we look at different variants. that give the False Positive rate $f$ of an $n+1$st message after $n$ messages have been inserted, e.g., one common variant yields $f=(1-(1-1/M)^{kn})^{k}$, where $M$ is the memory (number of bits in the BF) and $k$ is the number of hash functions, with each hash function mapping the message to one of the $M$ bits. This analysis, while useful to demonstrate efficacy of the BF, does not measure False Positives in a way that best serves practical network applications for two reasons:
Observability: First, consider the addition of an $n$th element. If the application maintains a count, $b$, of the number of bits set, the next new arrival being classified as a repeat (a False Positive) can be directly computed as $(b/M)^{k}$; the earlier analysis disregards the ability to inspect current BF state.
@@
A class of applications that would benefit from a measure of False Positive rate that takes into account both observability and average rate are those that employ what we call a Recycling Bloom Filter (RBF) [13, 20, 16]. These applications utilize BF’s across sufficiently long timescales, where new messages are continuously arriving, but are interspersed with repeats of previous messages. Since newly arriving messages almost always set bits in the BF, over time the number of bits will become excessive, driving up the False Positive rate, such that some bits must be cleared to keep False Positive rates at a reasonable level. While there are BF variants that do support deletion, their implementation introduces substantial overhead or requires knowledge of the specific elements destined for deletion [17, 10, 7, 2]. Such overhead and complexity is not suited for network applications where message arrivals have a “fading popularity” character, with repeat probability decreasing over time. For applications with this type of message arrival process, one can utilize an active metric that maintains some measure of BF fullness, such as the number of messages inserted, or the number of bits set in the BF. When this metric exceeds a threshold, the BF recycles: all bits are cleared, and the filling process begins anew (we say a new BF cycle is initiated). This process can continue indefinitely, providing a simple and effective means for maintaining a low maximal (and average) False Positive rate of arrivals.
In this paper, we develop recursive, quickly computable (on today’s conventional hardware) analytical models that accurately capture the average False Positive rates of Recycling BFs (RBFs). We show that using an active metric to gauge the recycle to maintain an average (as opposed to maximum) False Positive rate can offer a $30$% or more increase in amount of messages stored.
@@
  • We verify our models using discrete-event-driven simulation in §V, and, using our models, contrast performance of the RBF variants as functions of memory size and desired False Positive rates.
@@
Our findings show that the variant that recycles based on the number of bits set, while most complex to model to find the right parameters, is the best performing variant in that it can maintain a given False Positive rate while packing in the largest expected number of messages prior to recycling. In contrast, the variants that recycle based on message counts to achieve an average False Positive rate are easier to model and achieve only a slightly smaller expected number of messages prior to recycling. These average-case variants store around 30% more messages in expectation than the commonly used worst case variant. We also find that implementing a two-phase variant will closely match the one-phase as the False Positive rate to be met decreases. These results are important in that they point to users of BFs (and specifically RBFs) likely being overly conservative for their needs when using previous measures of False Positives.
@@ II Background
Markov Model internal variables
@@
False Positive Variants
$f_{w}$ False Positive rate using worst-case threshold (1)
$f_{o}$ False Positive rate using fixed $N$ oracle lower bound (2)
$f_{a}$ f.p. rate using fixed $N$ bit-setting msgs lower bound (3)
one-phase RBF (18)
$f_{\sigma}^{2}$ False Positive rate using $\sigma$ bound, two-phase RBF (26)
@@ II-A Bloom Filters
When a message arrives, the bits to which it hashes are checked before being set. If they are all already set, the message is assumed to be a repeat of a prior message. In the rare instance that a new message’s $k$ hashes point to bits already set, it is incorrectly classified as a repeat, i.e., a False Positive.
@@
Conventional implementations of hash functions permit several of the $k$ hash function mappings to collide and hash to the same bit (i.e., a pseudo-random sampling with replacement). A slightly more sophisticated hashing scheme can ensure the $k$ hashes for a message are distinct from one another (i.e., a pseudo-random sampling without replacement). We respectively refer to these two Bloom Filter variants as colliding and non-colliding, and we analyze both types. While we show how to incorporate these variants into our models, for values of $M$ and $k$ used in practice (typically $M>1,000$), the colliding/non-colliding distinction has an unnoticeable impact on False Positive rate (collisions are sufficiently rare in the colliding variant). Note that for either variant of hash functions, the independence of bits assigned to differing messages still holds. Hence, our results are generally presented using the colliding variant.
@@ II-B Traditional False Positive Analysis
where, for a given $M$ and $n$, $k\approx\frac{M}{n}\ln(2)$ will minimize the corresponding False Positive rate. Wikipedia provides a succinct derivation of this simple formula [21], as well as a formula for non-colliding hash functions (in terms of Stirling numbers). From (1), the False Positive rate approaches $1$ as the number of messages $n$ increases– an undesirable property for applications with no limit on the number of potential message insertions. We call using such a bound in practice to decide when to recycle the BF a worst case bound (hence $f_{w}$) because the recycling occurs as soon as a particular message’s False Positive rate exceeds $f_{w}$ (while the average over messages inserted will be lower). We refer to the lifetime of a BF between recycling events as a cycle of the BF.
@@ II-C Two-Phased Bloom Filter
An RBF, upon recycling, clears all memory with respect to previous arrivals. To mitigate the effects of this memory loss, a BF variant called a Two-Phase Bloom Filter has been used in various network applications, e.g., [13, 20].
The two-phase Bloom Filter can be thought of as two arrays with $M^{\prime}$ bits each, indexed $10,...,M^{\prime}-$ 1, along with $k$ independent hash functions that are shared across both arrays. Typically, $M^{\prime}=\frac{M}{2}$, where $M$ is what the application’s memory constraint would be for the standard (one phase) Bloom Filter.
The two-phase variant operates much like the standard recycling variant, with the addition that at any point in time, one Filter array is designated active and the other frozen. Incoming messages are hashed and added to the active filter. When the active filter is considered to be “full enough” (e.g., the number of bits set or messages stored exceeds some threshold), instead of clearing the bits of the active filter, we clear the bits of the frozen filter. The roles of the active and frozen filter are then switched; incoming messages are added to the previously frozen (now active) filter, and the status of the previously active (now frozen) filter remains unaffected. This process repeats indefinitely.
In the two-phase variant, an arriving message is hashed and its bits are compared separately within both the active and frozen filters. A match with either filter is interpreted as the message having been received previously (and having set the corresponding bits in the respective filter). While a two-phase RBF cannot reduce its false positive rate as effectively as its one-phase counterpart using identical total memory, its corresponding False Negative rates will often be significantly lower.
@@ II-D BF Thresholding Approaches
There are two basic approaches to deciding when to recycle a BF. One can count the number of bits $b$ that are set in the BF and recycle when $b$ exceeds a threshold $\sigma$. We call this approach a $\sigma$-bounded approach. Alternatively, one can bound the number of bit-setting messages, $N$, that can be admitted into the BF before recycling. [^2] Messages that don’t set bits are generally not counted because the user must assume they are repeat messages, and since repeat messages never set bits, the False Positive rate of the BF does not increase due to their repeated arrival. We call this approach an $N$-bounded approach. We note that one complication with an $N$-bounded approach is that when a message that arrives without setting any bits, a user cannot tell if it is a repeat message or a False Positive. A drawback of formulas for an $N$-bounded approach, such as (1), is that they count new messages regardless of whether or not new bits are set; since in practice new messages that do not set bits are assumed to be repeats, a user will underestimate the number of new messages received when False Positives occur, such that applications of a formula such as (1) to gauge when to recycle will be slightly off.
@@ II-E Averaging Variants
A BF is only of use when the message arrival process consists both of new and repeat messages. We wish to clarify a subtle point about what we are averaging over. When a message first arrives, if it sets additional bits, it is identified as new, and if not, it is classified as a repeat. Consider the $i$th new message that arrives, and let $\eta(i)$ represent the number of times the message arrives within a cycle, and let $F_{i}$ be an indicator that equals 1 if the $i$th message, upon first arrival, is a False Positive. We can count these $\eta(i)$ arrivals in three different ways.
@@
Which variant to use depends on the needs of the underlying application. For instance, a False Positive in a caching application that uses the BF to determine cache hits for requests would make the wrong assessment for an initial arrival of a request, but the mistake would be detected upon attempting to fetch the item from the cache, so the item would be cached and subsequent requests would not be false positives. Hence, count-first is the appropriate measure. Alternatively, a mechanism that wishes to count unique messages would have a single miscount for each message initially classified incorrectly as having been previously received, hence count-instance is the appropriate measure. Finally, in a setting where the BF is used to drop repeat requests, and where a penalty is paid per dropping of an unsatisfied request, count-each would be the appropriate measure.
We can prove that count-instance is the most stringent (largest) of False Positive rates of the three versions333count-first is straightforwardly smaller count-each.. The proof showing that count-instance results in the largest False Positive rate is omitted here but is included in §-A in the Appendix. This observation yields two benefits. First, if we compute the false positive rate for count-instance, we have upper bounds on the rates of the count-first and count-each variants. [^4] In fact, our analysis of count-instance is itself an upper-bound in an RBF for a subtle reason that we address in the extended version. Second, count-instance is not affected by underlying distribution of arriving messages: how often a particular message repeats is of no import, since we only decide its contribution to False Positive rate based on its first arrival, and any subsequent arrivals do not alter the bits in the BF, making its analysis identical regardless of this underlying distribution, whereas the other variants do depend on the underlying distribution.
@@ III $N$-bounded average False Positive rate
First, let us consider an oracle user, who, when inserting a message that sets no bits, can distinguish whether this message is a repeat, or a new message that failed to set bits. [^5] This oracle is a hypothetical user because the user can make such a distinction has no need for a BF. This oracle would insert exactly $N$ new messages, and the resulting average False Positive rate can be computed as a simple renewal process, yielding a count-instance oracle $N$-bounded average False Positive rate of:
@@
In contrast, a “real” user would assume that a new message that sets no bits is instead a repeat message, and would not include such a message in their count of $N$ messages. Hence, the oracle would recycle no later than an actual user, thereby ensuring a lower false positive rate. For this real user, we let $i$ iterate over the number of new messages that set bits, and define r.v. $R_{i}$ to equal the number of new messages that do not set bits and arrive between the $i-1$st and $i$th new messages that do set bits. Then for a single cycle of the RBF, the False Positive rate is $\sum_{i=1}^{N}R_{i}/\sum_{i=1}^{N}(R_{i}+1)$.
Note that for the messages that comprise a given $R_{i}$, the number of bits set in the BF remains constant (none of the $R_{i}$ messages are changing the number of bits set), so each such message has an equal likelihood, which we call $p_{i}$, of not setting additional bits. Note it follows that $P(R_{i}>j)={p_{i}}^{j+1}$, such that $E[R_{i}]=p_{i}(1-p_{i})^{-1}$. By renewal theory, we have the False Positive rate is thus equal to
@@
Finally, we show that $p_{i}$ is lower-bounded by $P(F_{i}=1)$. This follows from the fact that the $i$th new message that set bits is actually the $j$th arriving message with $j\geq i$, so $p_{i}=P(F_{j}=1)$ for some $j\geq i$. The lower bound then follows from the fact that $P(F_{i}=1)$ is increasing in $i$ (False Positive rates increase as more messages arrive), and that the worst-case analyses’ value for $N$ is over all messages, including those that set no bits. Hence, $p_{i}=P(F_{j}=1)\geq P(F_{i}=1)$ for $j\geq i$. Since we are underestimating each $p_{i}$, we are underestimating the likelihood of new arrivals not setting bits, so we are underestimating False Positive rates. Note that for low False Positive rates that are used in practice, $E[R_{i}]$ will tend to be very small, such that the bound is expected to be tight.
It will be useful for our purposes to define, for $f_{i}=P(F_{j}=1)$ and directly applying (1), a lower bound on the $N$-bounded False Positive rate of
@@ IV $\sigma$-bounded average False Positive rate
Figure 1: Partial Markov Chain diagram for Recycling Bloom Filter, showing the “forwards” transition behavior, for $k=2$ using colliding hash functions. States represent the number of RBF bits set to $1$ and $\tau_{k}{i,j}$ is the probability of transitioning from state $i$ to state $j$.
As a starting point towards developing a recursive analytical model that accurately captures the average False Positive rates for the $\sigma$-bounded model, we consider a process where messages sequentially arrive at the RBF. Once a message arrives, the $k$ hash functions are applied to it and the corresponding bits are set to $1$ (or remain $1$ if already set). In a departure from the $N$-bounded method of Bloom Filter recycling, we do not reset the RBF after $n$ unique message insertions. Instead, we have a predetermined bit capacity parameter $\sigma<M$. After the insertion of a message, let $b$ be the number of bits set to 1 in the RBF. We reset the RBF when $b>\sigma$. Our recursive Markov Model will allow us to determine the long-term average False Positive rate across all new arrivals for a given value of $\sigma$ (another departure from the traditional False Positive analysis, which only considers the “worst case” False Positive rate for the last arrival of a RBF reset cycle).